OpenStudy (anonymous):

If the derivative of f is given by f'(x) = (tan^2 x)/ (x^2+1) and f(1) = 1/2, then f(0) =?

myininaya (myininaya):

\[\int\limits_{}^{}\frac{\tan^2(x)}{x^2+1} dx\] we need to look at integrating this

OpenStudy (mr.math):

Are you sure it's \(\frac{\tan^2(x)}{x^2+1}\), I don't seem to be able to integrate that. It would be nicer if it was arctan instead of tan^2. :P

OpenStudy (turingtest):

I thought this was a FToC problem, which I kinda suck at...

OpenStudy (mr.math):

FToC?

OpenStudy (turingtest):

fundamental theorem of calculus

OpenStudy (mr.math):

Actually it is in a way or another.

OpenStudy (turingtest):

Right, so no need to integrate then, yes?

OpenStudy (mr.math):

We need to integrate.

OpenStudy (turingtest):

well wolfram quit

OpenStudy (mr.math):

We can find the value of f(0) numerically.

OpenStudy (anonymous):

yeah im sure its tan^2 x

OpenStudy (mr.math):

Call the definite integral \(\large \int_0^1 \frac{\tan^2(x)}{1+x^2}\) I, then: \[f(1)-f(0)=I \implies f(0)=f(1)-I\]. Now, try to evaluate I using any numerical method and solve substitute into the equation above.

OpenStudy (anonymous):

wat do u mean by any numerical method

OpenStudy (mr.math):

It's a method to finding approximate value for the integral. Try using Simpson's rule; it's easy and efficient.

OpenStudy (mr.math):

See here http://bit.ly/wg8SD3

OpenStudy (mr.math):

This is an approximate value for the integral, you can use it to find f(0).

OpenStudy (anonymous):

is it possible to get taht number without teh calcuator?

OpenStudy (mr.math):

What number?

OpenStudy (anonymous):

0.34456....which is the numeraical value of taht integral

OpenStudy (mr.math):

Yes, you can! Not necessarily this exact number. This approximation would probably require you to do 10's of iteration (I can't be sure), but you can get "an" approximation using Simpson's rule that's close to the exact value of the integral. http://en.wikipedia.org/wiki/Simpson's_rule

OpenStudy (anonymous):

lemme see if it works

OpenStudy (mr.math):

http://www.youtube.com/watch?v=ns3k-Lz7qWU

OpenStudy (anonymous):

here f(a) = f(0) but that is the integration of that (tan^2/x^2+1) which is vry hard 2 evaluate

OpenStudy (mr.math):

I don't understand what you mean.

OpenStudy (mr.math):

We're using a numerical method because we are unable to evaluate this integral, this is the whole point.

OpenStudy (anonymous):

yeah im trying to use the simpsons rule so wen i subsitute into it i get the expression 1/6 [ f(a)+ 4(1/2) + 1/2)] and istn f(a) taht integral

OpenStudy (anonymous):

i got the fromula from wikepedia

OpenStudy (mr.math):

No, the formula is considering f(x)=tan^2(x)/(1+x^2). make sense?

OpenStudy (anonymous):

not really

OpenStudy (mr.math):

Look closer at the formula :)

OpenStudy (mr.math):

You have f(x)=tan^2(x)/(1+x^2), a=0, and b=1. You have everything you need. For instance, f(0)=tan^2(0)/(1+0)=0.

OpenStudy (anonymous):

ohhhhhhhhh XD......

OpenStudy (mr.math):

I have to go in a few minutes, I think you will do fine.

OpenStudy (anonymous):

Yes i get it now! THANK YOU

OpenStudy (mr.math):

Glad to help! :)